# Bruno Bauwens 11 $0$-0002-6138-0591    Maria Marchenko 22

### v2

@@ Abstract
The even online Kolmogorov complexity of a string $x=x_{1}x_{2}\cdots x_{n}$ is the minimal length of a program that for all $i\leq n/2$, on input $x_{1}x_{3}\cdots x_{2i-1}$ outputs $x_{2i}$. The odd complexity is defined similarly. The sum of the odd and even complexities is called the dialogue complexity.
In [Bauwens, 2014] it is proven that for all $n$, there exist $n$-bit $x$ for which the dialogue complexity exceeds the Kolmogorov complexity by $n\log\frac{4}{3}+O(\log n)$. Let $\operatorname{C}^{s}(x)$ denote the Kolmogorov complexity with space bound $s$. Here, we prove that the space-bounded dialogue complexity with bound $s+6n+O(1)$ is at most $\operatorname{C}^{s}(x)+O(\log(sn))$, where $n=|x|$.
@@ 1 Introduction
The online Kolmogorov complexity is a generalization of conditional complexity, introduced in [7]. Given two lists of strings $x_{1},\ldots,x_{k}$ and $y_{1},\ldots,y_{k}$, it is defined by the minimal length of a program that for all $j\leq k$, on input $y_{1},\ldots,y_{j}$ prints $x_{j}$. This notion has important applications in the foundations of reinforcement learning [11, 12].111 An agent and an environment exchange messages. The AIXI-agent defined in [11, 12] models the environment using online a priori probability, which is a probabilistic version of complexity. The online coding theorem provides a close relation between the probabilistic and program complexity versions. This theorem is essential to prove the main result.
@@
In the limit for large $s$, this inequality reduces to additivity for unbounded complexity. A weaker result from [14, 15] is well known: if $s^{\prime}=2s+O(|xy|+\log s)$, then $\operatorname{C}^{s^{\prime}}(x)+\operatorname{C}^{s^{\prime}}(y|x)\leq\operatorname{C}^{s}(x,y)+O(\log s+\log\operatorname{C}^{s}(x,y))$. Because of the $O(\log s)$ in the right side, the result (*1) for unbounded complexity is not a corollary. The proof of this version follows the one for unbounded complexity.222 There is a small caveat: the decompresser needs to iterate over pairs $(x,y)$ with $\operatorname{C}^{s}(x,y)<k$, and hence, iterate through all halting programs that use at most $s$ bits of memory. To check whether a program halts, one needs an additional $s$ bits of memory. The main idea of the proof of theorem 1.2 is to do this more efficiently using Sipser’s trick from [18]: given a halting configuration of a machine, consider all computations that terminate in this state, and organize all states in a tree. Then do a DFS search starting in this halting configuration.
@@
Remark. It is not possible to drop the term $O(\log s)$, because otherwise, the limit for large $s$ would contradict theorem 1.1. Also, the proof gives a better precision of the big-O term at the right: $2d+4\log d+O(1)$ where $d=\operatorname{C}^{s}(s,\operatorname{C}^{s}(x),|x|)$.333 We believe it is possible to drop the term $O(d)$ when using prefix or monotone complexity, and still $s^{\prime}-s\leq O(|x|)$. However, the proof seems significantly harder. We hope to publish it in an extended version of this paper.
@@ 2 Definitions
  • One may think of the 2 stacks as a standard work tape, where 1 stack represents the part left of the computation head and other stack the right part. Note that to make this analogy work, one needs to think about a binary work tape with some “end marker” that can be moved.444 Popping an element from the stack corresponds to shifting all elements on one side of the tape, and this operation can be done if an endmarker is available. This definition allows for cleaner statements and proofs of space-bounded additivity (with $s^{\prime}=s+O(|xy|)$ without extra $O(\log s)$ term).
  • In proofs, one may use a more general type of machine, which is equipped with several standard binary work tapes that have end markers. Consider a machine with $k$ such tapes. Also, consider a computation on which at most $t_{1},\ldots,t_{k}$ bits of these tapes are used, and the total stack size is at most $s$. This computation is simulated on a machine of the above type using space $s+\sum_{i\leq k}t_{i}+2\lceil\log t_{i}\rceil$.
@@
Let $x,y\in\{0,1\}^{*}$. The conditional complexity is $\operatorname{C}^{s}(x|y)=\operatorname{C}^{s}(x{\leftarrow}y)$. If $y$ is empty, we write $\operatorname{C}^{s}(x)$. We use the prefix code $\bar{y}=0y_{1}\cdots 0y_{|y|}1$ and encode a tuple of strings $(t_{1},\ldots,t_{e})$ as $\bar{t}_{1}\bar{t}_{2}\cdots\bar{t}_{e-1}t_{e}$. The conditional complexity with tuples is defined using this encoding, for example $\operatorname{C}^{s}(y,z|x)=\operatorname{C}^{s}(\bar{y}z|x)$.
@@ 3 Even and odd online measures
– There are at most $12^{k}-$ 1 strings with complexity less than $k$, thus $Q(\text{empty})<1$.
@@
$\operatorname{C}^{s}(x)+1\geq-\log Q(x)=-\log\Big{(}\operatorname{P_{\textnormal{odd}}}(x)\,\operatorname{P_{\textnormal{ev}}}(x)\Big{)}\geq\operatorname{C}_{\textnormal{ev}}^{s^{\prime}}(x)+\operatorname{C}_{\textnormal{odd}}^{s^{\prime}}-O(\operatorname{C}^{s}(s,n,k)).$

### v4

@@ Abstract
The even online Kolmogorov complexity of a string $x=x_{1}x_{2}\cdots x_{n}$ is the minimal length of a program that for all $i\leq n/2$, on input $x_{1}x_{3}\cdots x_{2i-1}$ outputs $x_{2i}$. The odd complexity is defined similarly. The sum of the odd and even complexities is called the dialogue complexity. In [4] it is proven that for all $n$, there exist $n$-bit $x$ for which the dialogue complexity exceeds the Kolmogorov complexity by $n\log\frac{4}{3}+O(\log n)$. Let $\operatorname{C}^{s}(x)$ denote the Kolmogorov complexity with space bound $s$. Here, we prove that the space-bounded dialogue complexity with bound $s+6n+O(1)$ is at most $\operatorname{C}^{s}(x)+O(\log(sn))$, where $n=|x|$.
@@ 1 Introduction
The online Kolmogorov complexity is a generalization of conditional complexity, introduced in [7]. Given two lists of strings $x_{1},\ldots,x_{k}$ and $y_{1},\ldots,y_{k}$, it is defined by the minimal length of a program that for all $j\leq k$, on input $y_{1},\ldots,y_{j}$ prints $x_{j}$. This notion has important applications in the foundations of reinforcement learning [11, 12].111 An agent and an environment exchange messages. The AIXI-agent defined in [11, 12] models the environment using online a priori probability, which is a probabilistic version of complexity. The online coding theorem provides a close relation between the probabilistic and program complexity versions. A space bounded variant of this theorem is essential to prove the main result below.
@@
In the limit for large $s$, this inequality reduces to additivity for unbounded complexity. Only the weaker result from [14, 15] is well known: if $s^{\prime}=2s+O(|xy|+\log s)$, then $\operatorname{C}^{s^{\prime}}(x)+\operatorname{C}^{s^{\prime}}(y|x)\leq\operatorname{C}^{s}(x,y)+O(\log s+\log\operatorname{C}^{s}(x,y))$. The proof follows the one for unbounded complexity.222 There is a small caveat: the decompresser needs to iterate over pairs $(x,y)$ with $\operatorname{C}^{s}(x,y)<k$, and hence, iterate through all halting programs that use at most $s$ bits of memory. To check whether a program halts, one needs an additional $s$ bits of memory. The main idea of the proof of theorem 1.2 is to do this more efficiently using Sipser’s trick from [18]: given a halting configuration of a machine, consider all computations that terminate in this state, and organize all states in a tree. Then do a DFS search starting in this halting configuration. Because of the $O(\log s)$ in the right side, the result (*1) for unbounded complexity is not a corollary.
@@
Remark. It is not possible to drop the term $O(\log s)$, because otherwise, the limit for large $s$ would contradict theorem 1.1. Also, the proof gives a better precision of the big-O term at the right: $2d+4\log d+O(1)$ where $d=\operatorname{C}^{s}(s,\operatorname{C}^{s}(x),|x|)$.333 We believe it is possible to improve the $O(d)$ term to $O(\operatorname{C}_{s}(\operatorname{C}_{s}(s)))$ when using prefix or monotone complexity, and still $s^{\prime}-s\leq O(|x|)$. However, the proof seems significantly harder. We hope to publish it in an extended version of this paper.
@@ 2 Definitions
  • One may think of the 2 stacks as a standard work tape, where 1 stack represents the part at the left of the computation head and the other stack the right part. Note that to make this analogy work, one needs to think about a binary work tape with some “end marker” that can be moved.444 Popping an element from the stack corresponds to shifting all elements on one side of the tape, and this operation can be done if an endmarker is available. This definition allows for cleaner statements and proofs of space-bounded additivity (with $s^{\prime}=s+O(|xy|)$ without extra $O(\log s)$ term).
  • In proofs, one may use a more general type of machine, which is equipped with extra binary work tapes that have end markers. Consider a machine with $k$ such tapes. Also, consider a computation on which at most $t_{1},\ldots,t_{k}$ bits of these tapes are used, and the total stack size is at most $s$. This computation is simulated on a machine of the above type using space $s+\sum_{i\leq k}t_{i}+2\lceil\log t_{i}\rceil$.
@@
Note that even for $s=+\infty$, this definition is more general than in [7], because only $y_{j}$ is used to produce $x_{j}$, (not $y_{1},\ldots,y_{j}$).
@@
The conditional complexity of string $x$ given string $y$ is $\operatorname{C}^{s}(x|y)=\operatorname{C}^{s}(x{\leftarrow}y)$. If $y$ is empty, we write $\operatorname{C}^{s}(x)$. We use the prefix code $\bar{y}=0y_{1}\cdots 0y_{|y|}1$ and encode a tuple of strings $(t_{1},\ldots,t_{e})$ as $\bar{t}_{1}\bar{t}_{2}\cdots\bar{t}_{e-1}t_{e}$. The conditional complexity with tuples is defined using this encoding, thus $\operatorname{C}^{s}(y,z|x)=\operatorname{C}^{s}(\bar{y}z|x)$.
@@ 3 Even and odd online measures
– At most $12^{k}-$ 1 strings have complexity less than $k$, thus $Q(\text{empty})<1$.
@@
$\operatorname{C}^{s}(x)+1$ $\geq-\log Q(x)$
$=-\log\Big{(}\operatorname{P_{\textnormal{odd}}}(x)\,\operatorname{P_{\textnormal{ev}}}(x)\Big{)}$
$\geq\operatorname{C}_{\textnormal{ev}}^{s^{\prime}}(x)+\operatorname{C}_{\textnormal{odd}}^{s^{\prime}}-O(\operatorname{C}^{s}(s,n,k)).$
@@
Acknowledgements
The authors are grateful to an anonymous referee for suggesting a better title and careful comments.